Solve and show your work by completing the square.
x^2+3x-1=0
I have add 1 to both sides giving me x^2+3x=1 My example says half the 3 which would give me 1.5 and then square it but I am not getting an answer that my work program will except.

Answer by Alan3354(69443)

x^2+3x-1=0
I have add 1 to both sides giving me x^2+3x=1 My example says half the 3 which would give me 1.5 and then square it but I am not getting an answer that my work program will except [sic].
-----------------
x^2+3x = 1
x^2+3x + 2.25 = 1 + 2.25
(x + 3/2)^2 = 13/4
x + 3/2 = ±13/4
x = -3/2 ±13/4
-------------
Maybe it will accept fractions.